# DiscoTrack: A Multilingual LLM Benchmark for Discourse Tracking

### v2

@@ 1 Introduction
A growing number of benchmarks, such as GLUE (Wang et al., 2018) and CLUE (Xu et al., 2020), have been developed to evaluate LLMs across a wide range of phenomena, languages, and task types. However, the majority of existing benchmarks focus on extracting explicit information, often from individual sentences, and offer limited coverage of discourse-level phenomena that require reasoning across multiple sentences or larger contexts. Some recent efforts, such as Disco-Bench (Wang et al., 2023), have begun to address discourse-level evaluation. However, this benchmark is limited to only two languages and still focuses on QA and translation, rather than tracking discourse in terms of entities mentioned or implicit relations between utterances.
@@ 2 Related work
The most closely related effort to our work is Disco-Bench Wang et al. (2023), which specifically targets some discourse phenomena. However Disco-Bench is limited to Chinese and English, and targets discourse rather indirectly: for discourse understanding, it only uses three monolingual Chinese tasks: speaker identification, zero anaphora reconstruction (which is much less relevant for languages such as English) and classic reading comprehension questions. For generation it uses a mix of machine translation, text expansion and language modeling (text filling/completion), all of which benefit from discourse understanding, but do not probe discourse processing capabilities directly.
@@
This progression in task complexity is reflected in a variety of recent benchmarks. Early efforts such as GLUE and SuperGLUE (Wang et al., 2019, 2023) focused on localized classification tasks like natural language inference (NLI) and coreference linking. Later benchmarks like MMLU (Hendrycks et al., 2020), LM Evaluation Harness Gao et al. (2021) and BIG-Bench (Kazemi et al., 2025) expanded the scope to include general-purpose reasoning, knowledge retrieval, and problem-solving across domains. MT-bench (Bai et al., 2024) and LiveBench (White et al., 2024) have further shifted attention toward open-ended generation, assessing coherence and fluency in dialogue or multi-turn responses. Some efforts, such as AGIEval (Zhong et al., 2023), explicitly model human-like cognitive testing by evaluating performance on standardized exam questions. While many of these benchmarks and tasks involve reasoning over multiple sentences, they are not well suited to isolating discourse capabilities in the sense of computational discourse processing Stede (2012), i.e. with a focus on revealing implicit structure, priorities and unexpressed relationships that form the backbone of coherent texts.
@@ 3.1 Salience recognition
Salience in language can refer to a range of concepts (see Boswijk and Coler 2020), but for the present purpose, we define salience as the relative degree of prominence or importance that can be assigned to concepts in a document, which can be identified by a number of methods. In particular, our focus in this module of the benchmark is on Salient Entity Extraction (SEE, Lin and Zeldes 2024).
SEE datasets have taken a range of approaches for identifying the most relevant entities in context, including human annotation Dojchinovski et al. (2016), extraction from click-stream data Gamon et al. (2013), using hyperlinks and categories in Wikipedia data Wu et al. (2020) and automatically aligning entities extracted from documents using Named Entity Recognition (NER) with the same entities mentioned in summaries of those documents Dunietz and Gillick (2014). Unfortunately, almost all work on SEE has been conducted on English, meaning that in our benchmark this area is currently restricted to the evaluation of English language prompts only.
For our salience module, we employ a recently developed dataset annotated with graded salience scores, which range from 0–5 Lin and Zeldes (2025). The scores were computed by obtaining five summaries for each document in the freely available GUM corpus Zeldes (2017), which contains named and non-named entity and coreference annotations, and then manually aligning entities mentioned in the summaries to the document annotations. Entities mentioned in all five summaries receive the maximum score of 5, while entities mentioned in no summary score 0, using the reasoning that summaries will tend to mention the most salient entities in each document. The tasks in this module are:
@@ 3.2 Entity tracking
Entity tracking corresponds broadly to the well-studied task of coreference resolution Pradhan et al. (2014), in which spans of text that refer to the same entity in the world (e.g. ‘Mary’ and ‘her’ in Figure 1) are grouped into clusters. However, while structured coreference resolution can rely on pairwise linking each mention to just one antecedent to form clusters, LLMs reading a document holistically must implicitly track which expressions refer to the same entities in order to make correct inferences about discourse content. Rather than use structured coreference, which is a fairly unnatural task for text-driven LLMs, we employ a number of reasoning tasks more similar to a question-answering paradigm, which lend themselves to modeling with LLMs while forcing models to attend to entire documents.
Coreference information is available in a wide range of languages, and has recently been made available in the multilingual CorefUD dataset Nedoluzhko et al. (2022) in a uniform format, and we rely on this dataset to extract the subtask data for this module. The tasks are:
@@
The objective is to decide whether the two entities refer to the same thing in the world or not. In order to make the task more challenging than in previous benchmarks incorporating coreference, we pair true positives with ‘confusing’ negative examples, which we extract by running a recent multilingual neural coreference resolution system, MSCAW-coref Liu et al. (2024), on our documents, which is available as part of the Stanza NLP package Qi et al. (2020). We obtain negative pairs by selecting an anaphor predicted by the system for which we know a gold standard mention exists, and then select the highest probability incorrect antecedent considered by the system, again verifying that its mention exists in gold data. The filter for gold mentions ensures that no very easy cases are included which can be ruled out based on incorrect mention boundaries. A negative example is shown in 3.2.
@@ 3.3 Discourse relations
Multilingual identification of implicit and explicitly marked discourse relations such as cause or concession, and identifying connective words associated with such relations (e.g. because can mark cause as well as explanation) are crucial capabilities for the purpose of discourse understanding and explainable NLP Liu et al. (2023); Saeed et al. (2025). To test models’ ability to interpret discourse relations and their markers in text correctly, we assemble multilingual data for three text-based tasks, which we derive from the DISRPT shared task’s discourse relation classification and connective detection multilingual datasets Braud et al. (2024), as well as a number of additional datasets annotated according to the Penn Discourse Treebank (PDTB) framework Prasad et al. (2018). The tasks are as follows:
@@
In this task, the model is presented with two consecutive sentences (again, arg1 and arg2) that are connected by an implicit discourse relation, but for which no explicit connective appears in the text. The model is tasked with reconstructing the most appropriate connective that a human would use to connect the two sentences, making the relation explicit, as in 3.3. Negative examples are also provided, for which no connective is appropriate, and to which the correct response is ‘no implicit connective’, as in 3.3 (examples derived from GUM, Zeldes 2017).
@@ 3.4 Bridging inference
Bridging resolution is a task where bridging anaphora are automatically detected and resolved back to their associative antecedents. Thus far, bridging resolution has proven to be a difficult task, with SoTA systems achieving F-scores of < 0.5 on reference resources Kobayashi et al. (2023). Still, for an LLM to have full understanding of a discourse, an understanding of the implicit, associative relationships between bridging anaphora and their antecedents is essential.
For our benchmark evaluation, we use bridging annotations from the test set of the GUMBridge corpus Levine and Zeldes (2025), which each consist of a bridging anaphor and its associative antecedent. Each bridging pair in the test set is also categorized with a subtype annotation from the bridging subtype classification schema developed for GUMBridge (see the Appendix A for the subtypes).
@@ 4 Evaluation
To mitigate concerns about potential training data contamination, we note that several of our tasks are entirely new and unpublished or very recently published, including GUMBridge (Levine and Zeldes, 2025) and the new datasets and labels introduced in DISRPT 2025, which postdate training cutoffs on our tested models. The modest model scores below also suggest LLMs have probably not seen or memorized this data.
@@ 4.2 LLM performance
Overall, GPT-4o-mini and GPT-5 achieve the strongest results on most tasks, confirming their robustness across discourse-level evaluation, though as black box models, these results are not transparent or reproducible. The exception is bridgdet, where Aya-Expanse-8B few-shot outperforms other models. Following recent work on benchmark correlation and meta-analysis, (e.g. BenchBench, Zhang and Hardt 2024), we note that our model ranking largely agree with LMArena’s leaderboard (Chiang et al., 2024), where GPT-5 is followed by GPT-4o-mini, Aya-Expanse-8B, and Qwen-1.5-7B, though specific tasks and performance gap sizes do vary from overall LMArena rankings.
@@ Appendix A Prompts and Tag Sets
This appendix provides the full set of prompts and tag definitions used in our benchmark tasks. Section A first lists the exact prompts given to models for each task, followed by the tag sets used for classification. In total, we include 17 labels for core discourse relations, following the taxonomy of Bunt and Prasad (2016). For bridging anaphora, we adopt 11 fine-grained bridging types, as introduced by Levine and Zeldes (2025). These materials are included to ensure transparency and reproducibility.

### v3

@@ 1 Introduction
A growing number of benchmarks, such as GLUE (wang2018glue) and CLUE (xu2020cluecorpus2020), have been developed to evaluate LLMs across a wide range of phenomena, languages, and task types. However, the majority of existing benchmarks focus on extracting explicit information, often from individual sentences, and offer limited coverage of discourse-level phenomena that require reasoning across multiple sentences or larger contexts. Some recent efforts, such as Disco-Bench (wang2023discobench), have begun to address discourse-level evaluation. However, this benchmark is limited to only two languages and still focuses on QA and translation, rather than tracking discourse in terms of entities mentioned or implicit relations between utterances.
@@ 2 Related work
The most closely related effort to our work is Disco-Bench wang2023discobench, which specifically targets some discourse phenomena. However Disco-Bench is limited to Chinese and English, and targets discourse rather indirectly: for discourse understanding, it only uses three monolingual Chinese tasks: speaker identification, zero anaphora reconstruction (which is much less relevant for languages such as English) and classic reading comprehension questions. For generation it uses a mix of machine translation, text expansion and language modeling (text filling/completion), all of which benefit from discourse understanding, but do not probe discourse processing capabilities directly.
@@
This progression in task complexity is reflected in a variety of recent benchmarks. Early efforts such as GLUE and SuperGLUE (wang2019superglue; wang2023discobench) focused on localized classification tasks like natural language inference (NLI) and coreference linking. Later benchmarks like MMLU (hendrycks2020measuring), LM Evaluation Harness eval-harness and BIG-Bench (kazemi2025big) expanded the scope to include general-purpose reasoning, knowledge retrieval, and problem-solving across domains. MT-bench (bai2024mt) and LiveBench (white2024livebench) have further shifted attention toward open-ended generation, assessing coherence and fluency in dialogue or multi-turn responses. Some efforts, such as AGIEval (zhong2023agieval), explicitly model human-like cognitive testing by evaluating performance on standardized exam questions. While many of these benchmarks and tasks involve reasoning over multiple sentences, they are not well suited to isolating discourse capabilities in the sense of computational discourse processing Stede2012, i.e. with a focus on revealing implicit structure, priorities and unexpressed relationships that form the backbone of coherent texts.
@@ 3.1 Salience recognition
Salience in language can refer to a range of concepts (see BoswijkColer2020), but for the present purpose, we define salience as the relative degree of prominence or importance that can be assigned to concepts in a document, which can be identified by a number of methods. In particular, our focus in this module of the benchmark is on Salient Entity Extraction (SEE, lin-zeldes-2024-gumsley).
SEE datasets have taken a range of approaches for identifying the most relevant entities in context, including human annotation dojchinovski2016crowdsourced, extraction from click-stream data gamon2013identifying, using hyperlinks and categories in Wikipedia data wu2020wn and automatically aligning entities extracted from documents using Named Entity Recognition (NER) with the same entities mentioned in summaries of those documents dunietz2014new. Unfortunately, almost all work on SEE has been conducted on English, meaning that in our benchmark this area is currently restricted to the evaluation of English language prompts only.
For our salience module, we employ a recently developed dataset annotated with graded salience scores, which range from 0–5 lin2025gumsage. The scores were computed by obtaining five summaries for each document in the freely available GUM corpus zeldes2017gum, which contains named and non-named entity and coreference annotations, and then manually aligning entities mentioned in the summaries to the document annotations. Entities mentioned in all five summaries receive the maximum score of 5, while entities mentioned in no summary score 0, using the reasoning that summaries will tend to mention the most salient entities in each document. The tasks in this module are:
@@ 3.2 Entity tracking
Entity tracking corresponds broadly to the well-studied task of coreference resolution pradhan-etal-2014-scoring, in which spans of text that refer to the same entity in the world (e.g. ‘Mary’ and ‘her’ in Figure 1) are grouped into clusters. However, while structured coreference resolution can rely on pairwise linking each mention to just one antecedent to form clusters, LLMs reading a document holistically must implicitly track which expressions refer to the same entities in order to make correct inferences about discourse content. Rather than use structured coreference, which is a fairly unnatural task for text-driven LLMs, we employ a number of reasoning tasks more similar to a question-answering paradigm, which lend themselves to modeling with LLMs while forcing models to attend to entire documents.
Coreference information is available in a wide range of languages, and has recently been made available in the multilingual CorefUD dataset nedoluzhko-etal-2022-corefud in a uniform format, and we rely on this dataset to extract the subtask data for this module. The tasks are:
@@
The objective is to decide whether the two entities refer to the same thing in the world or not. In order to make the task more challenging than in previous benchmarks incorporating coreference, we pair true positives with ‘confusing’ negative examples, which we extract by running a recent multilingual neural coreference resolution system, MSCAW-coref liu-etal-2024-mscaw, on our documents, which is available as part of the Stanza NLP package qi-etal-2020-stanza. We obtain negative pairs by selecting an anaphor predicted by the system for which we know a gold standard mention exists, and then select the highest probability incorrect antecedent considered by the system, again verifying that its mention exists in gold data. The filter for gold mentions ensures that no very easy cases are included which can be ruled out based on incorrect mention boundaries. A negative example is shown in 3.2.
@@ 3.3 Discourse relations
Multilingual identification of implicit and explicitly marked discourse relations such as cause or concession, and identifying connective words associated with such relations (e.g. because can mark cause as well as explanation) are crucial capabilities for the purpose of discourse understanding and explainable NLP liu-etal-2023-hits; saeed-etal-2025-implicit. To test models’ ability to interpret discourse relations and their markers in text correctly, we assemble multilingual data for three text-based tasks, which we derive from the DISRPT shared task’s discourse relation classification and connective detection multilingual datasets braud-etal-2024-disrpt, as well as a number of additional datasets annotated according to the Penn Discourse Treebank (PDTB) framework prasad-etal-2018-discourse. The tasks are as follows:
@@
In this task, the model is presented with two consecutive sentences (again, arg1 and arg2) that are connected by an implicit discourse relation, but for which no explicit connective appears in the text. The model is tasked with reconstructing the most appropriate connective that a human would use to connect the two sentences, making the relation explicit, as in 3.3. Negative examples are also provided, for which no connective is appropriate, and to which the correct response is ‘no implicit connective’, as in 3.3 (examples derived from GUM, zeldes2017gum).
@@ 3.4 Bridging inference
Bridging resolution is a task where bridging anaphora are automatically detected and resolved back to their associative antecedents. Thus far, bridging resolution has proven to be a difficult task, with SoTA systems achieving F-scores of < 0.5 on reference resources kobayashi-etal-2023-pairspanbert. Still, for an LLM to have full understanding of a discourse, an understanding of the implicit, associative relationships between bridging anaphora and their antecedents is essential.
For our benchmark evaluation, we use bridging annotations from the test set of the GUMBridge corpus levine2025subjectivityannotationbridginganaphora, which each consist of a bridging anaphor and its associative antecedent. Each bridging pair in the test set is also categorized with a subtype annotation from the bridging subtype classification schema developed for GUMBridge (see the Appendix A for the subtypes).
@@ 4 Evaluation
To mitigate concerns about potential training data contamination, we note that several of our tasks are entirely new and unpublished or very recently published, including GUMBridge (levine2025subjectivityannotationbridginganaphora) and the new datasets and labels introduced in DISRPT 2025, which postdate training cutoffs on our tested models. The modest model scores below also suggest LLMs have probably not seen or memorized this data.
@@ 4.2 LLM performance
Overall, GPT-4o-mini and GPT-5 achieve the strongest results on most tasks, confirming their robustness across discourse-level evaluation, though as black box models, these results are not transparent or reproducible. The exception is bridgdet, where Aya-Expanse-8B few-shot outperforms other models. Following recent work on benchmark correlation and meta-analysis, (e.g. BenchBench, zhang2024inherent), we note that our model ranking largely agree with LMArena’s leaderboard (chiang2024chatbot), where GPT-5 is followed by GPT-4o-mini, Aya-Expanse-8B, and Qwen-1.5-7B, though specific tasks and performance gap sizes do vary from overall LMArena rankings.
@@ Appendix A Prompts and Tag Sets
This appendix provides the full set of prompts and tag definitions used in our benchmark tasks. Section A first lists the exact prompts given to models for each task, followed by the tag sets used for classification. In total, we include 17 labels for core discourse relations, following the taxonomy of bunt2016iso. For bridging anaphora, we adopt 11 fine-grained bridging types, as introduced by levine2025subjectivityannotationbridginganaphora. These materials are included to ensure transparency and reproducibility.